OpenStudy (anonymous):

Solve the following system of equations 2x - y + z = -3 2x + 2y + 3z = 2 3x - 3y - z = -4

OpenStudy (anonymous):

do u have a worksheet or book asighnment in alg 2???? because Im doing the same stuff with 3 equations and 3 variables?

OpenStudy (anonymous):

no, i'm doing school online.:) but yeah it is algebra ll

OpenStudy (anonymous):

set it up as a matrix and row reduce it

OpenStudy (anonymous):

im doing the same stuff. let me write the problem down. I might be able to help you, even though im kinda lost too. im writing the problem down

OpenStudy (anonymous):

these are the answers they gave me... but i dont get it:/ Select one: a. (1, 3, 2) b. (-1, 3, 2) c. (1, -3, 2) d. (1, 3, -2)

OpenStudy (anonymous):

how do you do that kimchyi?

OpenStudy (anonymous):

Have you learned about matrices? Or they want you to solve it by subsistiting /elimination?

OpenStudy (anonymous):

look this is all it says Solve the following system of equations. 2x - y + z = -3 2x + 2y + 3z = 2 3x - 3y - z = -4 (1 point) Select one: a. (1, 3, 2) b. (-1, 3, 2) c. (1, -3, 2) d. (1, 3, -2)

OpenStudy (anonymous):